A boy of mass 60 kg is standing over  a platform of mass 40 kg placed over a smooth horizontal surface. e throwsa stone of mass 1 kg with velocity v = 10 m/s at an angle of 45 ° w.r.t ground. Find the displacement of the platform (with boy) on the horizontal surface when the stone lands on the ground.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Sanchit

total mass of body and plateform = 100kg

 conservation of momentum in horizontal direction     100V + 1*10cos45 =0

                                                                               V=-1/10√2 m/sec 

 negative sign indicates in the reverse direction.

 time  of filght = 2*10 sin 45 /10

                    = √2

 so in this time plateform will move  = V*t

                                                    = 1/10√2 * √2

                                                    = .1 m answer

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
We are all IITians and here to help you in your IIT JEE & AIEEE preparation.

All the best.

Regards,
Askiitians Experts
Badiuddin